Meanwhile we managed to tag single alpha particles from an triple calibration source. Thereby we achived a spatial resolution of ~210ps timing resolution. After nearly one year of continuous operation we found some strange artefacts in the projections of the second unit. We found some strange sputter-like effects on the DDL and replaced it by a new, gold-coated one. After some tests with the new DDL the whole setup is running again. At the moment we are plannig a mounting and some modifications for the upcoming PRESPEC campaign, especially the accpeted UMBRELLA proposal, at the S4 degrader behind the FRS at GSI.

new lab and euroschool

Posted on September 15th, 2010 by Michael Pfeiffer

Recently came back from the 17th Euroschool on Exotic Beams with lots of ideas for improvment. Meanwhile we finally moved to the new lab and the whole setup is runnig again. We're trying to re-measure some of the latest results and/or to improve them. One can see the influence of the A/C in the new lab by watching at the stable vacuum (see below), not oscillating any more due to temperature shifts between night and day. Guess it's more an influence on the gauge meter itself :-)
image not available

new results and moving to a new lab

Posted on August 26th, 2010 by Michael Pfeiffer

We are back from our summer breaks. The data from the last experiments have been analysed and we managed to seperate the ions from a triple alpha source with good precision via TOF with a resolution of less then 400ps. Looking at the spatial resolution we still having issues concerning X/Y differences, but since the origin of that is known we're working of an improved prototype. Before all that can happen we have to move in a new lab, which is air conditioned. Hopefully it'll be done by the end of next week. Pictures will come :)

experiments keep running

Posted on July 20th, 2010 by Michael Pfeiffer

Meanwhile the new HV-supply is connected to CIMBI, so the online monitoring of the HV shows the current values now. We started doing some more experiments on the TOF resolution using a nearly un-coated Alpha-source and a mask with several small holes. Using a cut on those holes one can determine the TOF and its error for different flight-paths.

new high voltage supply

Posted on July 5th, 2010 by Michael Pfeiffer

The new CAEN N1471 HV-supply finally arrived. That's why the HV online monitor is up running again, even though it's just a test phase with the CIMBI not being connected.

the child has a name

Posted on June 28th, 2010 by Michael Pfeiffer

Meanwhile we began to systematically characterise the detectors response to the change of certain parameters. New NIM-based ultra fast TDCs are ordered, as well as matching TFAs and FAN-units to improve the whole acquisition system. Furthermore we are thinking about a circle shaped 150mm diameter version of the system.
We named the setup CIMBI, meaning Cologne Ion Monitor for Beam Imaging.

Below you can see a first result from the coincidence measurements using a mask showing the abbreviation of our institute. It is a projection on the position between both BPMs based on the measured trajectories of the alpha particles.
image not available

Pictures of the new setup for some coincidence tests are online.

Coincidence tests upcoming

Posted on April 19th, 2010 by Michael Pfeiffer

This week we'll hopefully start with some coincidence tests. Therefor we're putting a mask in front of the second detector, meaning the first one will only see 'noise' in MCA-mode whilst the second will see the mask projected by the ions. Taking both data in coincidence should 'cut out' the mask out of the noise of the first detector as function of the ions trajectories.

New pictures online

Posted on February 25th, 2010 by Michael Pfeiffer

Pictures of the renewed setup are online.

Redesign cycle nearly finished

Posted on February 9th, 2010 by Michael Pfeiffer

The redesign cycle is nearly finished. New implementations - besides the renewed table - are

  • a new online monitor for the high voltage,
  • a new solar cells trigger to fit the needs of
  • a new mask for the collimated beam,
  • a more stable mounting of the optical bench,
  • • • •
to mention the primary ones.

New Setup

Posted on Januar 19th, 2010 by Michael Pfeiffer

Meanwhile we've mounted the whole testbed on a completly renewed table. Hopefully we've reduced our noise-problems with the tables vibration a lot. Pictures and results are yet to come.

We started to modify the testbed for future experiments (e.g. LYCCA-DSSSD detectors as trigger, improved vaccum speed. etc.). Furthermore there are pictures of two new modifications available.

Complete Hamamatsu Setup

Posted on November 20th, 2009 by Michael Pfeiffer

Since the last post a lot happened. We got listmode data after our "first light" tests, we have two fully equipped and mounted Hamamatsu Setups running and we added really thin gold coated mylar foils for conversion (great thanks to Ganil at this point!). Furthermore we implemented a new source chamber and collimator (great thanks to our workshop!). Just to mention a few improvments! There will be some pictures of the 'new' setup, soon.

second "first light"

Posted on September 29th, 2009 by Michael Pfeiffer

Meanwhile the second detector is tested with high voltage as well. The first one was tested yesterday. Furthermore there is a new gallery of the laboratory online.

"first light"

Posted on September 28th, 2009 by Michael Pfeiffer

Today first tests with high voltage were performed with the first detector and we got promising signals. In beforehand only pulser signals (without high voltage being applied) were tested. Thus we got real signals from the 241Am source for the first time :)

About this Page / Page launched

Posted on September 22nd, 2009 by Michael Pfeiffer

This page is dedicated to the beam profile monitor being developed at the Institute of Nuclear Physics at the University of Cologne for the HiSpec/DeSpec project within the NuSTAR collaboration at FAIR.

It is intended to function as a kind of logbook for the progress of the project as well as a status monitor for the current test site. For the start, there is an online vacuum histogram and some pictures.